Circular No. 8669

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006aa, 2006ab, AND 2006ac Three apparent supernovae have been discovered on unfiltered KAIT

CCD images via the Lick Observatory Supernova Survey (cf. IAUC 8667), as reported by E. Lee, M. Baek, and W. Li:

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006aa Feb. 8.48 11^{h}53^{m}19.^{s}89 +20^{o}45′18′′.2 18.1 6′′.5 W, 12′′.2 N
2006ab Feb. 9.18 2 48 53.92 +53 02 20.8 18.0 9′′.2 E, 6′′.2 N
2006ac Feb. 9.47 12 41 44.86 +35 04 07.1 16.0 4′′.1 E, 21′′.7 N

Additional magnitudes of 2006aa in NGC 3947: Jan. 23.49 UT, [19.0; Feb. 9.50, 18.0. Additional magnitudes of 2006ab in PGC 10652: Jan. 7.20, [18.8; 24.16, [18.6; Feb. 10.18, 18.0. Additional magnitudes of 2006ac in NGC 4619, 2004 May 30.21, [19.0; 2006 Feb. 10.40, 15.9.

θ^{1} ORIONIS E

R. Costero, J. Echevarría, M. G. Richer, and A. Poveda, Universidad Nacional Autónoma de México, report their finding that the suspected variable star θ^{1} Ori E — the fifth brightest star in the Orion Trapezium, located ∼ 4′′ north of component A [object 1864 of Parenago 1954, Publ. Sternberg Astr. Inst. 25, 342; also ADS 4186 E, from Aitken’s 1932 Double Star Catalogue (Carnegie Inst. of Wash.), 1, 358; listed as NSV 2291 due to Walker, IBVS 1238] — is a double-lined spectroscopic binary. Echelle spectra (resolution ∼ 20000, range 380–680 nm), obtained on Jan. 8–17 with the 2.1-m telescope of the National Observatory at San Pedro Mártir, Baja California, show both components to be of spectral type F5. The Li I 670.8-nm line (equivalent width ∼ 0.14 nm) is clearly present in the spectra of both members. The estimated orbital period of the binary is 10.5 ± 1.2 days, and the maximum measured line separation is 154 ± 4 km/s. The spectrum of component E discussed by Herbig (1950, Ap.J. 111, 15) was probably contaminated by the light of component A; the present classification of component E is based on on the comparison of its spectrum with those of several standards of the Morgan and Keenan system, in the interval F0 V–G0 V, taken with the same instrument. These results, the suspected variability (0.4 mag), its being one of the brightest x-ray sources in the Trapezium (Schulz et al. 2003, Ap.J. 595, 365), and the position of the star well above the zero-age main sequence lead us to speculate that this object is an interesting candidate for its being an eclipsing system and/or a flare star. Further observations are encouraged.

2006 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green